# Distinguishing Power of 4-Legendrian Permutation Racks

### v1

@@ Abstract
We explore 4-Legendrian rack structures and the effectiveness of 4-Legendrian racks to distinguish Legendrian knots. We prove that permutation racks with a 4-Legendrain rack structure cannot distinguish sets of Legendrian knots with the same knot type, Thurston–Bennequin number, and rotation number.
@@ 1 Introduction
Joyce and Matveev both independently introduced quandles to the mathematical world in the 1980s [8, 13]. A few years later, Fenn and Rourke popularized a generalization of quandles, which they called “racks” [5]. These racks were first used to study framed knots and links, but it was quickly realized that every semi-framed non-split link embedded in a 3-manifold has a fundamental rack that classifies both the link and the 3-manifold.
In 2017, Kulkarni and Prathamesh [12] had the idea to use rack invariants to study Legendrian knots. In 2021, Ceniceros, Elhamdadi, and Nelson [1] generalized this into a more powerful structure called a Legendrian rack. In 2023, Kimura [10] and Karmakar, Saraf, and Singh [9] further independantly generalized these structures into the more powerful GL-rack, also called a generalized Legendrian rack or a bi-Legendrian rack. In 2024, Kimura [11] even further generalized GL-racks into 4-Legendrian racks, a potentially more powerful structure and the object of study for this paper.
In Section 2, we review common rack theory ideas and definitions and provide several examples of families of racks that we explore later in this paper.
In Section 3, we review the history of Legendrian racks up to this point, including many recent developments.
In Section 4, we define 4-Legendrian racks and discuss their classification as algebraic structures.
In Section 5, we review classical and rack-theoretic invariants of Legendrian knots.
In Section 6, we prove our main theorem, which is as follows.
@@
Kimura [10] and Cheng and He [2] both posed the open question of whether there exist 4-Legendrian racks that distinguish Legendrian knots sharing the same topological knot type and classical invariants. This question remains open in general. However, since permutation racks are the main examples of racks that are not quandles in the literature, our main theorem and a similar result of Kimura for 4-Legendrian quandles [*][Thm. 4.2.3]kimura-thesis settle this question in the negative for extremely large classes of 4-Legendrian racks.
@@ 2 Review of Racks
Definition 2.1 (Rack, [3]).
A rack, $(X,\triangleright)$, is a set X with binary operation $\triangleright:X\times X\to X$ satisfying:
@@
    R1 (Invertibility): For all y $\in$ X, the map $\beta_{y}:X\to X$ defined by $\beta_{y}(x)=x\triangleright y$ is invertible. We denote $\beta^{-1}_{y}(x)$ as $x\triangleright^{-1}y$.
@@
    R2 (Self-Distributivity): For all x, y, z $\in$ X, $(x\triangleright y)\triangleright z=(x\triangleright z)\triangleright(y\triangleright z)$.
R1 and R2 are known as the rack axioms. A rack is called a quandle when the binary operation is also idempotent, that is for all $x\in X,x\triangleright x=x.$ We say that $|X|$ is the order of $(X,\triangleright)$. Finite racks have been enumerated up to isomorphism for orders up to 13 and fully classified for orders up to 11 [18].
In the following, let $n\geq 0$ be a nonnegative integer.
@@
Let $t$ be an invertible variable, $s$ a non-invertible variable such that $s^{2}=s(1-t)$, and $X$ be a $\mathbb{Z}[t^{\pm},s]/(s^{2}-s(1-t))$-module. The rack with underlying set $X$ and rack operation $x\triangleright y\coloneq tx+sy$ is called a $(t,s)$-rack.
@@
All (t, 1-t)-racks are Alexander quandles. When $s\neq 1-t$, the resulting rack is not a quandle as $x\triangleright x\neq x$.
@@ 2.1 Rack automorphisms
It can be shown that $\pi$ is a rack automorphism that commutes with all other rack homomorphisms; that is, $\pi$ lies in the center of the category of racks. Moreover, the inverse map is given by $\pi^{-1}(x)=x\triangleright^{-1}x$. See, for example, [15, Sec. 2.2].
@@
3 History of Legendrian Racks
Although rack theory [5] and Legendrian knot theory [4] were popularized concurrently in the 1990s, it was not until 2017 when Kulkarni and Prathamesh [12] introduced the first rack-theoretic invariants of Legendrian knots. They called these invariants $n$-Legendrian racks, and they used these invariants to distinguish infinitely many Legendrian unknots [12, Main Thm. 2]. In 2021, Ceniceros, Elhamdadi, and Nelson [1] generalized $n$-Legendrian racks by introducing Legendrian racks, which are 4-Legendrian racks where $u_{l}=u_{r}=d_{l}=d_{r}$. They used Legendrian racks to distinguish certain Legendrian trefoils and connected sums of Legendrian trefoils [1, Sec. 5].
Further generalizing Legendrian racks, Kimura [10] and Karmakar, Saraf, and Singh [9] independently introduced GL-racks (also called generalized Legendrian racks or bi-Legendrian racks) in 2023. GL-racks are 4-Legendrian racks where $u_{l}=u_{r}$. They distinguish infinitely many Legendrian unknots and Legendrian trefoils; see [10, Thm. 4.1] and [9, Thms. 4.7 and 4.8]. In 2025, Cheng and He [2, Thm. 1.1] further showed that GL-racks distinguish Legendrian knots at least up to the absolute values of their classical invariants.
However, none of the above examples of Legendrian knots answer the question of whether rack invariants can distinguish Legendrian knots with the same topological knot type and classical invariants. In fact, in his 2024 PhD thesis introducing 4-Legendrian racks, Kimura [11, Sec. 4.2] found several examples of nonequivalent Legendrian knots sharing these invariants that cannot be distinguished using 4-Legendrian racks. He also showed that 4-Legendrian quandles cannot distinguish any such Legendrian knots [11, Thm. 4.2.3]. Our main theorem (Theorem 6.3) extends this result to all 4-Legendrian permutation racks. It is worth noting that while Kimura’s result uses a theorem from contact topology, our approach is purely algebro-combinatorial.
On the algebraic side of things, the first author [15] in 2025 presented a simplified but equivalent characterization of GL-racks. This led to a group-theoretic classification of GL-structures, answering a question posed by Karmakar, Saraf, and Singh in a previous version of [9]. Combining [15, Thm. 5.6] and [16, Thm. 10.1] yields the surprising result that the categories of racks, Legendrian racks, and GL-quandles are isomorphic. Moreover, certain involutory GL-racks have connections to symmetric racks (also called racks with good involutions) [*][Thm. 10.1]Ta_2025goodinvolutionsconjugationsubquandles, which are used to distinguish surface-knots in $\mathbb{R}^{4}$.
@@ 3 History of Legendrian Racks
Remark 4.3.
Remark 4.4.
Since $U_{X}$ is nonempty, every rack $(X,\triangleright)$ can be equipped with at least one 4-Legendrian structure. A similar statement holds for GL-racks but not for Legendrian racks.
@@
Example 4.7 (Cf. [10, Ex. 3.6]).
Let $G$ be a group, and let $g,h\in Z(G)$. Let $u_{l},u_{r}\colon G\to G$ be the multiplication maps $x\mapsto gx$ and $x\mapsto hx$. Then $(\operatorname{Conj}(G),u_{l},u_{r})$ is a 4-Legendrian quandle.
@@
In his PhD thesis introducing 4-Legendrian racks, Kimura [11, Sec. 4.2] defined a 4-Legendrian rack as a sextuple $(X,\triangleright,u_{l},u_{r},d_{l},d_{r})$ in which $(X,\triangleright)$ is a rack and $u_{l},u_{r},d_{l},d_{r}\colon X\to X$ are functions that satisfy the following eight axioms for all $x,y\in X$:
$d_{l}u_{r}=u_{r}d_{l}$ $=d_{r}u_{l}=u_{l}d_{r},$ (4.1)
@@
$x\triangleright d_{l}(y)$ $=x\triangleright y=x\triangleright d_{r}(y).$
@@
With some work, one can show that Kimura’s definition of 4-Legendrian racks is equivalent to our definition; $d_{l}$ and $d_{r}$ are respectively determined entirely by $u_{r}$ and $u_{l}$ as well as $\triangleright$:
@@
In fact, the forgetful functor defined by $(X,\triangleright,u_{l},u_{r},d_{l},d_{r})\mapsto(X,\triangleright,u_{l},u_{r})$ is an isomorphism of categories. The proof of this fact is similar to the analogous result for GL-racks [15, Prop. 3.12]; we omit the details here.
In light of this equivalence, we use our definition while studying algebraic properties of 4-Legendrian racks, and we use Kimura’s definition while studying Legendrian knot invariants.
@@
Given a 4-Legendrian permutation rack as in Example 4.6, we have
@@ 4.1 Classification of 4-Legendrian Racks
Let $(u_{l},u_{r})$ and $(v_{l},v_{r})$ be 4-Legendrian structures on $(X,\triangleright)$. By definition, the 4-Legendrian racks $(X,\triangleright,u_{l},u_{r})$ and $(X,\triangleright,v_{l},v_{r})$ are isomorphic if and only if there exists a rack automorphism $\varphi\in\operatorname{Aut}(X)$ such that $v_{l}=\varphi u_{l}\varphi^{-1}$ and $v_{r}=\varphi u_{r}\varphi^{-1}$. Equivalently, the pairs $(u_{l},u_{r})$ and $(v_{l},v_{r})$ are simultaneously conjugate in $\operatorname{Aut}R$; that is, the 4-Legendrian structures lie in the same orbit of $U_{X}\times U_{X}$ under the diagonal conjugation action of $\operatorname{Aut}(X)$. This action exists because $U_{X}$ is a normal subgroup of $\operatorname{Aut}(X)$. To summarize, we have just shown the following.
Proposition 4.10 (Cf. [15, Thm. 4.1]).
@@
Let $n\geq 0$ be a nonnegative integer, and let $T_{n}$ be the trivial quandle of order $n$. Then $U_{X}=\operatorname{Aut}(T_{n})=S_{n}$, so isomorphism classes of $4$-Legendrian racks with underlying rack $T_{n}$ correspond to orbits of $S_{n}\times S_{n}$ under the diagonal action of $S_{n}$ by conjugation. These orbits are counted in OEIS sequence A110143 [7]; see [17]. (We verified this example for all $n\leq 6$ using the computer search described below.)
@@
Even if $(X,\triangleright,u_{l})\cong(X,\triangleright,v_{l})$ and $(X,\triangleright,u_{r})\cong(X,\triangleright,v_{r})$ as GL-racks, the 4-Legendrian racks $(X,\triangleright,u_{l},u_{r})$ and $(X,\triangleright,v_{l},v_{r})$ are not necessarily isomorphic. For example, let $T_{3}$ be the trivial quandle of order $3$, and let $u_{l}=u_{r}=v_{l}=(2,3)$ and $v_{r}=(1,3)$ in cycle notation.
@@
Given a 4-Legendrian rack $(X,\triangleright,u_{l},u_{r})$, it is usually not isomorphic to $(X,\triangleright,u_{r},u_{l})$. For example, let $u_{l}$ be any nonidentity GL-structure on a rack $(X,\triangleright)$, and let $u_{r}\coloneq\operatorname{id}_{X}$.
Using Proposition 4.10, we implemented a GAP [6] program that classifies all $4$-Legendrian racks of a given order $n\leq 11$ up to isomorphism. This program uses the classification of racks up to order $11$ from [18], which is where the $n\leq 11$ bound comes from.
We were able to complete the search for all $n\leq 6$; our code and data is available in a GitHub repository [14]. Table 1 enumerates our data. The table also includes an enumeration of involutory 4-Legendrian racks; this is motivated by the connection between involutory GL-racks and symmetric racks shown in [16, Sec. 10].

### v2

@@ Abstract
We study 4-Legendrian racks and their effectiveness at distinguishing Legendrian knots. We prove that permutation racks with 4-Legendrian rack structures cannot distinguish Legendrian knots that share the same knot type, Thurston–Bennequin number, and rotation number. However, they also recover these three classical invariants.
@@ 1 Introduction
In 2023 and 2025, Kimura [Kim23] and Cheng and He [CH25] posed the question of whether certain algebraic objects called 4-Legendrian racks can distinguish Legendrian knots with the same topological knot type, Thurston–Bennequin number, and rotation number. In this work, we provide a negative answer for all 4-Legendrian permutation racks.
@@
Permutation racks with 4-Legendrian structures cannot distinguish isotopy classes of Legendrian knots with the same classical invariants.
Since permutation racks are the main examples of racks that are not quandles in the literature, Theorem 1.1 and a similar result of Kimura for 4-Legendrian quandles [*][Thm. 4.2.3]kimura-thesis settle this question in the negative for very large classes of 4-Legendrian racks.
On the other hand, we also show that the Legendrian knot invariants provided by 4-Legendrian racks are at least as powerful as the classical invariants.
Theorem 1.2 (Corollary 5.3).
The fundamental 4-Legendrian rack $F(L)$ recovers the classical invariants of $L$.
Theorem 1.2 can be viewed as a stronger version of the main theorem in a paper of Cheng and He [CH25] from 2025. To prove their result, Cheng and He computed a normal form for words in certain colorings of Legendrian knots by GL-racks, which are a special class of 4-Legendrian racks. By adapting this method to 4-Legendrian racks more generally, we obtain a key lemma (Lemma 5.2) used to prove both Theorem 1.1 and Theorem 1.2.
Kimura introduced 4-Legendrian racks in his 2024 PhD thesis [Kim24] by equipping algebraic structures called racks with extra structure. Racks are generalizations of algebraic structures called quandles, which Joyce [Joy82] and Matveev [Mat84] independently introduced in the 1980s to construct invariants of knots and links. In 1992, Fenn and Rourke [FR92] introduced racks to study framed knots and links. It was quickly realized that every semi-framed non-split link embedded in a 3-manifold has a fundamental rack that classifies both the link and the 3-manifold.
Although rack theory and Legendrian knot theory [Etn05] were popularized concurrently in the 1990s, it was not until 2017 when Kulkarni and Prathamesh [KP17] introduced the first rack-theoretic invariants of Legendrian knots. They called these invariants $n$-Legendrian racks, and they used these invariants to distinguish infinitely many Legendrian unknots [KP17, Main Thm. 2]. [^1] Despite the name, 4-Legendrian racks are not a special class of $n$-Legendrian racks; in fact, the former vastly generalizes the latter. In 2021, Ceniceros, Elhamdadi, and Nelson [CEN21] generalized $n$-Legendrian racks by introducing Legendrian racks, which are 4-Legendrian racks where $u_{l}=u_{r}=d_{l}=d_{r}$; cf. (3.1). They used Legendrian racks to distinguish certain Legendrian trefoils and connected sums of Legendrian trefoils [CEN21, Sec. 5].
Further generalizing Legendrian racks, Kimura [Kim23] and Karmakar, Saraf, and Singh [KSSar] independently introduced GL-racks (also called generalized Legendrian racks or bi-Legendrian racks) in 2023. GL-racks are 4-Legendrian racks where $u_{l}=u_{r}$ (and, hence, $d_{l}=d_{r}$; cf. (3.1)). They distinguish infinitely many Legendrian trefoils and even more Legendrian unknots than $n$-Legendrian racks can; see [Kim23, Thm. 4.1] and [KSSar, Thms. 4.7 and 4.8]. In 2025, Cheng and He [CH25, Thm. 1.1] further showed that GL-racks distinguish Legendrian knots at least up to the absolute values of their classical invariants.
However, none of the above examples of Legendrian knots answer the question of whether rack invariants can distinguish Legendrian knots with the same topological knot type and classical invariants. Kimura [Kim23] and Cheng and He [CH25] both posed this question. In his PhD thesis, Kimura [Kim24, Sec. 4.2] discovered several examples of nonequivalent Legendrian knots sharing these invariants that cannot be distinguished using 4-Legendrian racks. He also showed that 4-Legendrian quandles cannot distinguish any such Legendrian knots [Kim24, Thm. 4.2.3]. Our main theorem (Theorem 1.1) extends this result to all 4-Legendrian permutation racks. We note that while Kimura’s result uses a theorem from contact topology, our approach is purely algebro-combinatorial.
On the algebraic side of things, the first author [Ta25a] studied GL-racks from group-theoretic, categorical, and universal-algebraic perspectives in 2025. In particular, a simplified but equivalent definition of GL-racks led to a group-theoretic characterization of GL-structures, answering a question posed by Karmakar, Saraf, and Singh in a previous version of [KSSar]. Combining [Ta25a, Thm. 5.6] and [Ta25b, Thm. 10.1] yields the surprising result that the categories of racks, Legendrian racks, and GL-quandles are isomorphic. Moreover, certain involutory GL-racks have connections to symmetric racks (also called racks with good involutions) [*][Thm. 10.1]Ta_2025goodinvolutionsconjugationsubquandles, which are used to distinguish surface-knots in $\mathbb{R}^{4}$. These connections formalized an observation that Karmakar, Saraf, and Singh made in a previous version of [KSSar].
This paper is organized as follows. In Section 2, we review common rack theory ideas and definitions and provide several examples of families of racks that we explore later in this paper. In Section 3, we define 4-Legendrian racks and discuss their classification as algebraic structures. In Section 4, we review classical and rack-theoretic invariants of Legendrian knots. In Section 5, we prove Theorems 1.1 and 1.2. In Section 6, we propose areas for future research.
@@ 2 Review of Racks
Definition 2.1 (Rack, [EN15]).
A rack $(X,\triangleright)$ is a set X with binary operation $\triangleright\colon X\times X\to X$ satisfying the following two axioms:
@@
    R1 (Invertibility): For all y $\in$ X, the map $\beta_{y}\colon X\to X$ defined by $\beta_{y}(x)=x\triangleright y$ is invertible. We denote $\beta^{-1}_{y}(x)$ by $x\triangleright^{-1}y$.
@@
    R2 (Self-distributivity): For all $x,y,z\in X$, we have $(x\triangleright y)\triangleright z=(x\triangleright z)\triangleright(y\triangleright z)$.
A rack is called a quandle if the binary operation $\triangleright$ is idempotent, that is, $x\triangleright x=x$ for all $x\in X$. We say that $|X|$ is the order of $(X,\triangleright)$. Finite racks have been enumerated up to isomorphism for orders up to 13 and fully classified for orders up to 11 [VY19].
In the following, let $n\geq 1$ be a positive integer.
@@
Let $X$ be a module over the quotient ring $\mathbb{Z}[t^{\pm},s]/(s^{2}-s(1-t))$. The rack with underlying set $X$ and rack operation $x\triangleright y\coloneq tx+sy$ is called a $(t,s)$-rack.
@@
All (t,1t)(t,1-t)-racks are Alexander quandles. When $s\neq 1-t$, the resulting rack is not a quandle as $x\triangleright x\neq x$.
@@ 2.1 Rack automorphisms
One can show that $\pi$ is a rack automorphism; the inverse map $\pi^{-1}$ is given by $x\mapsto x\triangleright^{-1}x$. One can also show that $\pi$ and $\pi^{-1}$ commute with all other rack homomorphisms, implying that $\pi$ and $\pi^{-1}$ lie in the center of the category of racks. See, for example, [Ta25a, Sec. 2.2].
@@
Example 3.5 (Cf. [Kim23, Ex. 3.6]).
Let $G$ be a group, and let $g,h\in Z(G)$. Let $u_{l},u_{r}\colon G\to G$ be the multiplication maps $x\mapsto gx$ and $x\mapsto hx$, respectively. Then $(\operatorname{Conj}(G),u_{l},u_{r})$ is a 4-Legendrian quandle.
@@
For all racks $(X,\triangleright)$, the maps $\operatorname{id}_{X}$, $\pi$, and $\pi^{-1}$ are GL-structures, though they are not necessarily all distinct. Thus, every rack can be equipped with at least one 4-Legendrian structure; a similar statement holds for GL-racks but not for Legendrian racks.
We comment on the relationship between our definition of 4-Legendrian racks and the original definition. In his PhD thesis introducing 4-Legendrian racks, Kimura [Kim24, Sec. 4.2] defined a 4-Legendrian rack as a sextuple $(X,\triangleright,u_{l},u_{r},d_{l},d_{r})$ in which $(X,\triangleright)$ is a rack and $u_{l},u_{r},d_{l},d_{r}\colon X\to X$ are functions that satisfy the following eight axioms for all $x,y\in X$:
$d_{l}u_{r}=u_{r}d_{l}$ $=d_{r}u_{l}=u_{l}d_{r},$
@@
$x\triangleright u_{l}(y)$ $=x\triangleright y=x\triangleright u_{r}(y).$
@@
With some work, one can show that Kimura’s definition of 4-Legendrian racks is equivalent to our definition. Namely, $d_{l}$ and $d_{r}$ are respectively determined entirely by $u_{r}$ and $u_{l}$ as well as $\triangleright$:
@@
In fact, the forgetful functor defined by $(X,\triangleright,u_{l},u_{r},d_{l},d_{r})\mapsto(X,\triangleright,u_{l},u_{r})$ is an isomorphism of categories. This fact is proven similarly to the analogous result for GL-racks [Ta25a, Prop. 3.12]; we omit the details here.
In light of this equivalence, we use our definition to study algebraic properties of 4-Legendrian racks, and we use Kimura’s definition to study Legendrian knot invariants.
@@
Given a 4-Legendrian permutation rack as in Example 3.4, combining (3.1) with Example 2.16 yields
@@ 3.1 Classification of 4-Legendrian Racks
Let $(u_{l},u_{r})$ and $(v_{l},v_{r})$ be 4-Legendrian structures on $(X,\triangleright)$. By definition, the 4-Legendrian racks $(X,\triangleright,u_{l},u_{r})$ and $(X,\triangleright,v_{l},v_{r})$ are isomorphic if and only if there exists a rack automorphism $\varphi\in\operatorname{Aut}(X)$ such that $v_{l}=\varphi u_{l}\varphi^{-1}$ and $v_{r}=\varphi u_{r}\varphi^{-1}$. Equivalently, the pairs $(u_{l},u_{r})$ and $(v_{l},v_{r})$ are simultaneously conjugate in $\operatorname{Aut}(X)$; that is, the 4-Legendrian structures lie in the same orbit of $U_{X}\times U_{X}$ under the diagonal conjugation action of $\operatorname{Aut}(X)$. This action exists because $U_{X}$ is a normal subgroup of $\operatorname{Aut}(X)$. To summarize, we have just observed the following.
Proposition 3.9 (Cf. [Ta25a, Thm. 4.1]).
@@
Let $n\geq 0$ be a nonnegative integer, and let $T_{n}$ be the trivial quandle of order $n$. Then $U_{X}=\operatorname{Aut}(T_{n})=S_{n}$, so isomorphism classes of $4$-Legendrian racks with underlying rack $T_{n}$ correspond to orbits of $S_{n}\times S_{n}$ under the diagonal action of $S_{n}$ by conjugation. These orbits are counted in OEIS sequence A110143 [Han05]; see [Tra]. (We verified this example for all $n\leq 6$ using the computer search described below.)
@@
Even if we have GL-rack isomorphisms $(X,\triangleright,u_{l})\cong(X,\triangleright,v_{l})$ and $(X,\triangleright,u_{r})\cong(X,\triangleright,v_{r})$, the 4-Legendrian racks $(X,\triangleright,u_{l},u_{r})$ and $(X,\triangleright,v_{l},v_{r})$ are not necessarily isomorphic. For example, let $T_{3}$ be the trivial quandle of order $3$, and let $u_{l}=u_{r}=v_{l}=(2,3)$ and $v_{r}=(1,3)$ in cycle notation.
@@
Given a 4-Legendrian rack $(X,\triangleright,u_{l},u_{r})$, it is not necessarily isomorphic to $(X,\triangleright,u_{r},u_{l})$. For example, let $(X,\triangleright)$ be any rack such that $U_{X}$ contains a nonidentity element $\varphi\neq\operatorname{id}_{X}$, and let $(u_{l},u_{r})\coloneq(\varphi,\operatorname{id}_{X})$.
Using Proposition 3.9, we implemented a GAP [24] program that classifies all $4$-Legendrian racks of a given order $n\leq 11$ up to isomorphism. This program uses the classification of racks up to order $11$ from [VY19], which is where the $n\leq 11$ bound comes from.
We were able to complete the search for all $n\leq 6$; our code and data is available in a GitHub repository [Ta25]. Table 1 enumerates our data. The table also enumerates 4-Legendrian involutory racks and 4-Legendrian kei; this is motivated by the connections between involutory GL-racks and symmetric racks shown in [Ta25b, Sec. 10].